How many different numbers can be formed with the digits 2,4,6,8,0 each being divisible by ten, repetition of the digits being not allowed?
Answer by checkley71(8403)

(4*3*2*1)+(4*3*2)+(4*3)+4=24+24+12+4=64
5 DIGIT NUMBERS +4 DIGIT NUMBERS + 3 DIGIT NUMBERS +2 DIGIT NUMBERS ALL ENDING IN 0.
EXAMPLES:
86240
2640
640
40